OpenStudy (anonymous):

solve by the substitution method 5x-5y=-20 7x+46=y the solution is ( ) type and ordered pair or there is no solution

OpenStudy (anonymous):

5x-5(7x+46)=-20 5x-35x-230=-20 -30x=230-20 -30x=210 x=-7 y=7x+46 y=7(-7)+46 y=-49+46 y=-3

OpenStudy (anonymous):

replace y in the first equation by 7x+46 to get \[5x-5(7x-46)=-20\] and solve for x

OpenStudy (anonymous):

what pat said!

OpenStudy (anonymous):

the ordered pair is (-7,-3)

OpenStudy (anonymous):

satellite? what do you mean?

OpenStudy (anonymous):

i mean you are right!

OpenStudy (anonymous):

ah, i see. i thought, it is the other way around..lol

OpenStudy (anonymous):

so no reason for me to continue to write it out.

OpenStudy (anonymous):

i am sorry if i am too fast to answer.

OpenStudy (anonymous):

thanks pat18

OpenStudy (anonymous):

you are welcome^_^

OpenStudy (anonymous):

Pat18 can you help me with this one it is a little different solve by the substitution method 4m+n=21 m-2n=12

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

4m+n=21 m-2n=12------> m=2n+12 4(2n+12)+n=21 8n+48+n=21 9n=21-48 9n=-27 n=-3 m=2n+12 m=2(-3)+12 m=-6+12 m=6

OpenStudy (anonymous):

thanks again pat 18 you're very smart this stuff is new to me i have not had math in years

OpenStudy (anonymous):

so pat what is the ordered pair or is there no solution

OpenStudy (anonymous):

the ordered pair is (-3,6) for (n,m)

OpenStudy (anonymous):

Pat 18 could you recheck the system tells me that is not correct

OpenStudy (anonymous):

this is correct. i got the correct answer.

OpenStudy (anonymous):

ok maybe I made a typo this is the correct way to type and ordered pair (-3,6) correct

OpenStudy (anonymous):

yeah, that is what i got.

OpenStudy (anonymous):

Pat18 the system say to begin by solving either equation for either variable, substitute the resultiing expression into the other equation and solve to find the value of one variable. Use back-substitution to find the value of the other variable. there is no solution if both variables are eliminated and a false equation reamains.

OpenStudy (anonymous):

yeah, i used the substitution.i substituted the value of y in the equations.

OpenStudy (anonymous):

so there is no solution